Circular No. 8598

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005du, 2005dv, 2005dw, 2005dx, 2005dy, 2005dz Six apparent supernovae discoveries have been reported from unfiltered

CCD images. SN 2005du was found by the ‘Nearby Supernova Factory’ group via G. Aldering (cf. CBET 214); 2005dv by P. Dainese (Marghera, Italy) and A. Dimai (Cortina, Italy) in the course of the CROSS program (cf. IAUC 7373); 2005dw and 2005dx by N. Ponticello, M. Baek, J. Burket, and W. Li (via LOSS/KAIT; cf. IAUC 8594); and 2005dy by A. Sehgal, J. Newton, and T. Puckett with the 0.50-m Sehgal telescope (Osoyoos, BC) in the course of the Puckett Observatory Supernova Search (cf. IAUC 8580). SN 2005dz was found independently by Puckett and A. Pelloni with a 0.60-m reflector (whence the tabulated data below) and by LOSS.

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005du Aug. 8.5 0^{h}09^{m}34.^{s}7 +12^{o}45′21′′.4 19.7
2005dv Sept. 4.83 13 41 04.45 +67 40 20.3 15.0 8′′ W
2005dw Aug. 30.29 22 13 31.83 −27 33 19.6 16.7 6′′.4 E, 11′′.1 N
2005dx Sept. 9.49 4 04 28.85 −17 54 27.5 18.6 12′′.3 W, 10′′.8 S
2005dy Sept. 7.36 4 28 38.72 +70 20 49.0 19.0 5′′.0 W, 1′′.3 S
2005dz Sept. 10.21 23 38 38.33 + 5 26 17.3 18  24′′ E, 16′′ N

Additional approximate magnitudes from the respective observing groups: SN 2005dv in NGC 5283, June 27 UT, [19.0; Aug. 5, [18.5; Sept. 6.79, 14.8. SN 2005dw in MCG –05-52-49, Aug. 12.28, [18.5; Sept. 9.29, 16.9. SN 2005dx in MCG –03-11-9, Feb. 23.17, [19.5; Sept. 11.48, 18.7. SN 2005dy in UGC 3042, July 27, [20.5; Aug. 29, [20.5; Sept. 11.35, 18.7. SN 2005dz in UGC 12717, July 27, [19.8 (Puckett); 29, [19.8 (Puckett); Sept. 2.30, [19.0 (KAIT); 11.31, 18.1 (KAIT). SN 2005du was confirmed photometrically on Aug. 11.5 with the Lick 1-m Nickel reflector and was found to be a type-II supernova via spectroscopic observations with the 5-m Palomar reflector by A. Gal-Yam, D. Sand, and D. C. Leonard on Aug. 14. Also, nothing is visible at the location of SN 2005dv on Palomar Sky Survey red, blue, and infrared images. Li provides position end figures 38^{s}.28, 17′′.1 (and offsets 23′′.8 east, 15′′.0 north of the center of UGC 12717) for SN 2005dz.

F. J. D. Serduke, D. S. Wong, K. Shimasaki, and A. V. Filippenko, University of California, Berkeley, report that inspection of CCD spectra (range 330–1000 nm), obtained on Sept. 11 UT with the Shane 3-m reflector at Lick Observatory, shows that SN 2005dx and SN 2005dw are relatively young type-II supernovae. In both cases, broad Hα emission is the strongest feature.

2005 September 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green